# Improving U-Net Confidence on TEM Image Data with L2-Regularization, Transfer Learning, and Deep Fine-Tuning

### v1

@@ 4 Encoder Fine-Tuning
Figure 7: Metrics obtained from 5-fold cross-validation for fine-tuning optimization. The horizontal axis refers to which blocks of EfficientNetB7 (see Fig. 3) are unfrozen. “Frozen” refers to no fine-tuning and “Random” refers to no pre-training. Errors bars correspond to ±$\pm 1\sigma$
@@
It has been shown that the performance of U-Net can be improved by using an EfficientNetB7 encoder pre-trained on ImageNet, with all weights unfrozen. However, it remains unclear whether performance gains stem from the EfficientNet architecture itself, or from the benefits of pre-training and fine-tuning. To address this question, model performance was studied for various degrees of encoder fine-tuning and L2-regularization. Starting from a fully frozen encoder with ImageNet weights, the blocks of EfficientNet were progressively unfrozen until all were trainable. To isolate the effect of pre-training, an additional condition was tested in which all weights were fully randomized and models were trained end-to-end. The evaluation results across these fine-tuning levels are summarized in Fig. 7, where the horizontal axes represent the degree of fine-tuning and the vertical axes are the metrics and their values. For each condition, 5-fold cross-validation was again used to obtain robust performance estimates. Additionally, three levels of L2-regularization (with $\lambda$=0, 1e-4, 1e-3) were applied to assess its potential influence on the observed trends.
@@
According to Fig. 7, the first key observation is that when no fine-tuning is applied (i.e., left-most on the horizontal axes), models are skeptical with high precision, low recall, and low self-confidence. While increasing regularization leads to modest improvements, model self-confidence rises significantly only after unfreezing deeper encoder blocks. Second, all metrics roughly plateau once block 5 is unfrozen, with Fig. 7e showing the highest grain detection rate when $\lambda$ = 1e-3. Third, when trained from scratch (i.e., right-most on the horizontal axes), however, both certainty and abundance drop substantially. Furthermore, grain detection rate decreases by 20%. These results indicate pre-training is necessary for maximizing performance with EfficientNet encoders, and that fine-tuning the deepest layers leads to maximum confidence.
@@
The trends seen in Fig. 7 provide some insight into the abilities of EfficientNet to adapt its pre-trained weights for a new task and dataset. The steep rise in performance as more blocks become trainable implies that the most important dissimilarity between ImageNet and the TEM dataset used in this work is the fine image structures. That is, most improvement comes from developing shallow layers responsible for extracting localized features. Although less significant quantitatively, developing deep layers which extract global features are crucial for building model confidence. This is most likely due to TEM images being much more homogeneous than natural scene images like those in ImageNet, and thus more suitable weights for global feature extraction must be learned [12].
@@ 5 Conclusion
In this work, the performance of U-Net on a TEM image dataset of nanocrystalline grain morphologies has been improved by using a pre-trained EfficientNetB7 encoder with deep fine-tuning and L2-regularization. Two novel metrics, prediction certainty and abundance, have been introduced to provide a qualitative evaluation without regard a ground truth. A third metric, grain detection rate, was also used to provide a practical and holistic evaluation of model performance.
Replacing the encoder with EfficientNetB7 gave models excess capacity and redundancy, and by using L2-regularization, they obtained substantially higher prediction confidence. In total, 57% more grains were detected over the unregularized baseline U-Net architecture. To explain these benefits, it was proposed that regularization reduces model self-uncertainty by preventing overfitting of arbitrary human errors, but is only useful when a large pre-trained backbone is used.
A deeper dive into the importance of encoder pre-training was conducted as well, where it was found that model self-confidence is only acquired by fine-tuning the deepest layers which have been pre-trained on ImageNet. Some insight into the dissimilarity between domain-specific and natural scene images used for training was discussed and it was concluded that there exists important differences at each spatial level during encoding. By fine-tuning all weights, responsible for extracting local and global features, model performance reaches its maximum.
@@ Acknowledgments
All authors were supported by the U.S. Department of Energy’s Distinguished Early Career Award, project DE-NE0009426. Special thanks to Arthur Motta for providing us with the TEM images used in this work, as well as Zefeng Yu, Aiping Chen, Yogesh Sharma, and Wei-Ying Chen for data creation (UO2 sample fabrication, TEM imaging, and analysis).
Code and Data Availability
The code used in this work is located at https://github.com/psu-rdmap/unet-compare and can be used to train models on our dataset, or any other dataset with the allowable format. The data and analysis code, except for the model files due to size constraints, can be found on Zenodo with DOI 10.5281/zenodo.15626325.

### v2

@@ 4 Encoder Fine-Tuning
Figure 7: Metrics obtained from 5-fold cross-validation for fine-tuning optimization. The horizontal axis refers to which blocks of EfficientNetB7 (see Fig. 2) are unfrozen. “Frozen” refers to no fine-tuning and “Random” refers to no pre-training. Errors bars correspond to $\pm 1\sigma$
@@
It has been shown that the performance of U-Net can be improved by using an EfficientNetB7 encoder pre-trained on ImageNet, with all weights unfrozen. However, it remains unclear whether the performance gains stem from the EfficientNet architecture itself, or from pre-training and fine-tuning. To address this question, model performance was studied for various degrees of encoder fine-tuning and L2-regularization. Starting from a fully frozen encoder with ImageNet weights, the blocks of EfficientNet were progressively unfrozen until all were trainable. To isolate the effect of pre-training, an additional condition was tested in which all weights were fully randomized and models were trained end-to-end. The evaluation results across these fine-tuning levels are summarized in Fig. 7, where the horizontal axes represent the degree of fine-tuning and the vertical axes show the metrics and their values. For each condition, 5-fold cross-validation was again used to obtain robust performance estimates. Additionally, three levels of L2 regularization (with $\lambda$=0, 1e-4, 1e-3) were applied to assess its potential influence on the observed trends.
@@
According to Fig. 7, the first key observation is that when no fine-tuning is applied (i.e., left-most on the horizontal axes), models are skeptical with high precision, low recall, and low self-confidence. While increasing regularization leads to modest improvements, model self-confidence rises significantly only after unfreezing deeper encoder blocks. Second, all metrics roughly plateau once block 5 is unfrozen, with Fig. 7e showing the highest detected defect count when $\lambda$=1e-3. Third, when trained from scratch (i.e., right-most on the horizontal axes), both certainty and abundance drop substantially. Furthermore, detected defect count decreases by 25% for $\lambda$=1e-3. These results indicate pre-training is necessary for maximizing performance with EfficientNet encoders, and that fine-tuning deeper layers up to block 5 leads to maximum self-confidence.
@@
The trends seen in Fig. 7 provide some insight into the abilities of EfficientNet to adapt its pre-trained weights for a new task and dataset. The steep rise in F1-score when the first few blocks (i.e., blocks 1-2) become trainable implies that the most important dissimilarity between ImageNet and the TEM dataset used in this work is the fine image structures. That is, most improvement in F1-score comes from developing shallow layers responsible for extracting localized features. However, developing deep layers up to block 5, which extract more global features, are crucial for building model self-confidence. This is most likely due to TEM images being much more homogeneous than natural scene images in ImageNet, and thus unique weights for global feature identification must be learned [13]. However, very abstract and compressed features discovered in the deepest layers remain relatively consistent between ImageNet and TEM datasets.
@@ 5 Conclusion
In this work, the performance of U-Net on a TEM image dataset of nanocrystalline grain morphologies has been improved by using a pre-trained EfficientNetB7 encoder with deep fine-tuning and L2-regularization. Two novel metrics, prediction certainty and abundance, have been introduced to provide a qualitative evaluation independent of a ground truth. A third metric, detected defect count, was also used to provide a practical and holistic evaluation of model performance.
Replacing the encoder with EfficientNetB7 gave models excess capacity and redundancy, and by using L2-regularization, they obtained substantially higher prediction confidence. In total, 64% more grains were detected over the unregularized baseline U-Net architecture. To explain these benefits, it was proposed that regularization enhances model self-confidence by preventing overfitting of arbitrary human errors, but is only useful when a large pre-trained encoder is used.
A deeper dive into the importance of encoder pre-training was conducted as well, where it was found that model self-confidence is only acquired by fine-tuning deep layers which have been pre-trained on ImageNet, perhaps due to the dissimilarity between TEM and natural scene images at each spatial level during encoding. By fine-tuning blocks 1-5, responsible for extracting local and global features, model performance reaches its maximum.
Code and Data Availability
The code developed for this work can be found on GitHub at https://github.com/psu-rdmap/unet-compare. Training results, accompanying analysis code, and the supplementary materials can be found on ScholarSphere at https://scholarsphere.psu.edu/resources/b80356d7-6485-40aa-841d-8f598c4ee9e2.
@@ Acknowledgments
This work was supported by the Office of Nuclear Energy of U.S. Department of Energy under Contract No.DE-NE0009426. We thank Dr. Zefeng Yu and Prof. Arthur Motta of Pennsylvania State University for providing the original TEM images.